# Finding radio transients with anomaly detection and active learning based on volunteer classifications

### v1

@@ 3.4 Feature space and clustering
An obvious question to ask is how well each feature set separates transients (anomalies) from the rest of the distribution. For a perfect feature set, we would want all transients to be separated cleanly from all non-anomalies. To inspect this we make use of a Uniform Manifold Approximation and Projection (UMAP, McInnes et al., 2018), a technique for dimensionality reduction. UMAP attempts to project data from its N-dimensional space (where N is 45 and 40 for the feets and wavelet features respectively and 2 for $\eta$ and $V$) down into 2-D space for visualisation, whilst preserving local structure.
As a baseline, the feature space of ($\eta,V$) is shown in Figure 2. This figure demonstrates that many transients and variables are inliers, as found in A23. The presence of systematics in some fields (e.g. MAXI J1820) can be seen, along with the difficulty in separating our transients (anomalies) from all background sources. This was one of the main motivations for testing other feature sets for anomaly detection.
@@
Figure 3: The UMAP of our feets feature space, colour-coded by observational field. We see that clustering is mainly a function of observational field, aligning with the heterogeneous nature of the dataset. We note that the axes are unlabelled as these numbers are arbitrary combinations of many features and have no physical meaning. The inset is a demonstration of how the data of one particular field (GX339) is distributed, as mentioned in the text.
@@ 4 Anomaly detection and active learning
Having calculated sets of features for our light curves, we move on to the core anomaly detection step of Astronomaly. To do this we pass our standardised features into two commonly used anomaly detection algorithms, detailed below, as implemented in the scikit-learn package (Pedregosa et al., 2011). Anomaly detection is typically done by each algorithm having a definition of what is normal and assigning subjects that are far in parameter space from normal as highly anomalous, without any user labels or training. Once anomalousness has been ranked by a given algorithm, Astronomaly scales this to a human reading of 0 to 5, where 5 is most anomalous. By its nature, unsupervised learning has no pre-defined metric to calculate success. However, as noted earlier, we have votes from citizen scientists for all of our light curves from BfS:MKT, which in turn have been verified by us. Throughout this work we will assume that these verified volunteer votes are an absolute ground truth against which to compare our anomaly detection methods. We can then use the positions of our citizen-science verified transients in our anomaly score-ordered list to calculate metrics of success - the more transients that are ranked as highly anomalous, the better. It is important to point out that Astronomaly never decides whether or not something is anomalous and simply orders the entire list by anomaly score. The use of information from our volunteers is what allows us to then calculate metrics of success such as recall and purity (see section 4.2). A contrasting method for cases without such citizen science data might be to define some threshold in an anomaly score-ordered list above which things are said to be anomalies, providing the binary classes on which to calculate performance.
@@ 4.1 Anomaly detection algorithms
The first of the two anomaly detection methods used herein is the Local Outlier Factor (LOF, Breuniq et al., 2000). The LOF is a density-based method that measures the deviation of the density of a given sample with respect to some user-defined number of nearest neighbours, $k$. We set $k=100$ throughout this work, approximately 1 per cent of our sample. Values of $k$ between 20 (the default) and 200 were tested, producing changes in performance at the sub per cent level. As noted earlier, without labelled data it would be difficult to optimise this (and other) hyperparameters. The LOF is local in that the resultant anomaly score depends on how isolated the object is with respect to the surrounding neighbourhood. More specifically, the LOF of datum A is a measure of the average local density of all k neighbours divided by A’s own local density. One advantage of this local approach is that it might pick up anomalies missed by global density measurements. That is, a point only a small distance form a very dense cluster might be an anomaly, whereas a point in a sparse cluster might still have the same density as its neighbours and so not be an outlier despite its low absolute density value. For more on the distinction between global and local anomalies, see the discussion in Rogers et al. (2024).
@@
4.2 Anomaly Recall and Purity
In order to assess the performance of our models we can ask out of all the anomalies that exist, what fraction did we find in the top n sources? This quantity is known as the recall (or completeness) and is only calculable with ground truth labels for an entire dataset, as without this it is impossible to know how many anomalies are missed (false negatives). Recall is defined as
@@ 4.2 Anomaly Recall and Purity
where for us the denominator is our total number of transients (anomalies) in the data, i.e. 168. This recall value, whilst simple, corresponds well with preferable performance consisting of many more anomalies high in the list for the user to see quickly. We can calculate the recall $R$ as a function of n until we reach the end of our dataset ($n=8874)$, by which time we will have recovered all sources in the list. We define $R_{10}$ and $R_{30}$ as the recall of transients in the top 10 and 30 per cent of our list respectively, which we will use as quantitative measures of performance throughout this work. We stress again that it is impossible to calculate this recall without labels for all transients, as without this we would not known how many transients fall below a given threshold i.e. the rate of false negatives. As such, looking at the top 10 or 30 per cent of the anomaly score-ordered lists is somewhat arbitrary and depends on, for example, the resources available for follow up or further citizen science classification.
We can also ask what fraction of the top $n$ sources are transients. This is known as the precision or purity, defined as
@@
This can be thought of as a measure of the hit-rate of our anomalies. Namely, how many transients (anomalies) would there be in the top 10 or 30 per cent of the data? We note that, by definition, as $n$ approaches the total length of our dataset (8874 sources), the purity will tend towards the aggregate rate of transients in our data i.e. $\sim 2$ per cent. Purity, unlike recall, only depends on the labels of sources above some cut-off point set by the user and so is calculable even for a largely unlabelled dataset.
@@ 4.3 Active Learning
In this work we actively label the top $\sim$ 2 per cent of each feature-model pair, which takes of order 10 minutes by an individual (AA), applying a consistent methodology throughout. That is, examples of clear variability or known transients were ranked highly ($U=4$ or 5), whilst low-significance light curves, or those showing clear false positive behaviour - e.g. where one epoch shows large changes in flux due to a different calibrator or worse image quality - were ranked at 1 or 0. However, it is worth noting that this labelling is inherently subjective and dependent on a user’s end goal. An instrument scientist interested in flux calibration would score sources differently to an astronomer interested in stellar radio flares. Active learning attempts to quantify this and balance an algorithmic definition of anomalous with a user-dependent idea about what is an interesting anomaly. Finally we note that the regressor used to model user interest is by default a RF but that using a Gaussian process provides an improvement on this for some datasets, which will be incorporated into future versions of Astronomaly (Walmsley et al., 2022).
@@ 5.1 Baseline Anomaly Detection
Figure 6: Upper panel: The recall curves for both anomaly detection algorithms as applied to three sets of features derived from our radio light curves. The data are ordered according to the anomaly score from each algorithm, where for example $x=0.3$ corresponds to the most anomalous thirty per cent of the dataset. The dotted line in each corresponds to $y=x$ i.e. random chance. Lower panel: A combined plot of all six recall lines for better comparison between feature sets. Dashed lines make comparisons (see text) for recall rates at 10 and 30 per cent of the dataset. The values in the legend correspond to the area under each recall curve (AUC).
@@
The recall curves for both anomaly detection algorithms applied to all three feature sets can be seen in Figure 6. These are normalised on both axes i.e. x=0.1 corresponds to the top 10 per cent most anomalous sources and y=0.5 corresponds to half of the bona fide 168 variables found by volunteers. Better performance corresponds to curves with higher gradient at low $n$, appearing pushed up into the upper left corner. It can be clearly seen that ($\eta,V$) show poor performance on this dataset, essentially matching random chance, in line with A23’s findings. As noted earlier, this will be due to a combination of applying a different methodology to the intended use of $\eta$ and $V$ and the inherent biases of these two statistics - e.g. $\eta$ is correlated with signal-to-noise squared (see equation 2). Furthermore, our ground truth data consist of sources chosen irrespective of their ($\eta,V$) values. Contrastingly, we see that our main feature sets are able to recover a majority of our volunteer-verified transients (anomalies) in a minority of the data. This is the first demonstration of anomaly detection as applied to radio time series.
@@
When comparing performance between anomaly detection models (i.e. for the same feature set), we see minimal variation and qualitatively similar behaviour between recall curves. For the feets and wavelet features we see ‘broken power-law’ type behaviour, with an approximately constant gradient initially, followed by a knee and then worse performance. Table 2 lists all $R_{10}$ and $R_{30}$ values for model-feature set pairs. The largest discrepancies between models occur at the ‘top’ of our list, with $R_{10}$ values varying by 0.1 for feets features and 0.14 for wavelets. However, $R_{30}$ values differ from 0.03 to 0.07, indicating that model choice most affects the highly anomalous sources in our dataset. It is worth noting that several iterations of each feature-model pair were tested, with performance never varying by more than $\sim$1 per cent. As a result, all values are rounded to this precision.
@@
Differences in feature set have a much greater effect than changing the anomaly detection method - for the same model $R_{10}$ changes by factors of $\sim$3.5 between best- and worst- performing feature sets. Similarly, $R_{30}$ for the baseline parameters reaches only $\sim$0.32, whilst the wavelet features can recover up to 91 per cent of volunteer-verified transients. This underscores the importance of using features that best describe a given dataset.
@@
We can also calculate an aggregate recall performance by simply integrating the area under the curve (AUC) of each model-feature set’s recall curve. The corresponding AUC values for each model-feature set instance are listed in the legend of Figure 6. As already noted, differences between feature sets are greater than differences between anomaly detection models. These AUC values show that the wavelet-IF combination has the best recall performance at an AUC of 0.83. However, this is a summary over all the data and does not consider the actual finding of transients in a live setting. By this we mean that one must consider how many anomalies a human will be able to find in a reasonable amount of time. As discussed above and can be seen in the lower panel of Figure 6, if a user can only inspect the top 10 per cent of the dataset, then the feets-LOF analysis performs the best, with $R_{10}=0.52$. However, if there are resources available to inspect 30 per cent of the dataset (e.g. computation, citizen scientists, a team of experts) then the LOF performance on wavelet features recalls 91 per cent of transients, 7 per cent better than the next-best feature-model pair ($R_{30}=0.84)$.
feets Wavelets $(\eta,V)$
LOF IF LOF IF LOF IF
$R_{10}$ 0.52, 0.52 0.42, 0.45 0.24, 0.32 0.38, 0.42 0.14, 0.16 0.12, 0.11
$R_{30}$ 0.74, 0.73 0.78, 0.78 0.91, 0.91 0.84, 0.84 0.34, 0.32 0.31, 0.30
$P_{10}$ 0.11, 0.11 0.09, 0.10 0.05, 0.06 0.07, 0.08 0.03, 0.03 0.02, 0.02
$P_{30}$ 0.05, 0.05 0.05, 0.05 0.06, 0.06 0.05, 0.05 0.02, 0.02 0.02, 0.02
@@
Table 2: Specific recall ($R$) and purity ($P$) values, rounded to two decimal places, for each model-feature set pair, measured after ten and thirty per cent of the most anomalous sources have been ‘seen’. Each pair of values listed are the default, active iteration of that feature-model instance.
@@
The purity values for all feature-model pairs can be found in Table 2, again looking at 10 and 30 per cent of our anomaly score-ordered lists. For the feets and wavelet feature sets, purity is always at least a factor of 2 better than the corresponding baseline performance. In general, the same trends hold as with our recall values - differences between feature sets are greater than differences between anomaly detection models. Likewise, we can see that the feets features produce the best results when looking at the top 10 per cent of the data, with a hit-rate of over 1 in 10, compared to the baseline of 1 in 50. The results 30 per cent of the way through the anomaly score-ordereds, as seen with the recall values, show that the wavelet-LOF instance performs the best, but only marginally. Nevertheless, we can see that for either feature set, as applied to both of our anomaly detection algorithms, the hit-rate of anomalies in the data is always a factor of 2 or greater when compared to our baseline. This directly translates into the number of sources that would need to be sifted out before finding a new transient.
@@
In summary, the choice of model used makes a small difference, though LOF is marginally preferred. When considering feature sets, the baseline features perform the worst, recovering transients almost uniformly through the anomaly score-ordereds. The feets features perform the best at the beginning of each anomaly score-ordered, recovering 52% of transients in 10% of the data and increasing the purity above random by a factor of 5. However, the wavelet features perform better when considering a larger fraction of each anomaly score-ordered. Using the wavelet features over 90 per cent of anomalies are recovered in 30 per cent of the data, at a purity level that is three times higher than the baseline features.

### v2

@@ 3.4 Feature space and clustering
An obvious question to ask is how well the feature sets and baseline feature pair separate transients (anomalies) from the rest of the distribution. For a perfect feature set, we would want all transients to be separated cleanly from all non-anomalies. To inspect this we make use of a Uniform Manifold Approximation and Projection (UMAP, McInnes et al., 2018), a technique for dimensionality reduction. Briefly, the UMAP algorithm works by building a topological representation of data in its N-dimensional space (where N is 45 and 40 for the feets and wavelet features respectively). Once this has been constructed, the UMAP algorithm works to optimize a low dimensional (e.g. 2-D) representation of this topological space as measured by cross entropy. Datapoints that are similar in the N-D space should then remain nearby in the 2-D representation, which can then be used for visualisation or further downstream tasks.
As a baseline, the feature space of ($\eta,V$) is already in 2-D so there is no need to reduce its dimensionality. This feature space is shown in Figure 2. This figure demonstrates that many transients and variables are inliers, as found in A23. The presence of systematics in some fields (e.g. MAXI J1820) can be seen, along with the difficulty in separating our transients (anomalies) from all background sources. This was one of the main motivations for testing the subsequent feature sets for anomaly detection.
@@
Figure 3: The UMAP of our feets feature space, colour-coded by observational field. We see that clustering is mainly a function of observational field, aligning with the heterogeneous nature of the dataset. The axes are unlabelled as these numbers are arbitrary combinations of many features and have no physical meaning. The inset is a demonstration of how the data of one particular field (GX339) is distributed, as mentioned in the text.
@@ 4 Anomaly detection and active learning
Having calculated features for our light curves, we move on to the core anomaly detection step of Astronomaly. To do this we pass our standardised features into two commonly used anomaly detection algorithms, detailed below, as implemented in the scikit-learn package (Pedregosa et al., 2011). Anomaly detection is typically done by each algorithm having a definition of what is normal and assigning subjects that are far in parameter space from normal as highly anomalous, without any user labels or training. Once the degree to which something is anomalous has been ranked by a given algorithm, Astronomaly scales this to a human reading of 0 to 5, where 5 is most anomalous. The output algorithm scores are linearly rescaled such that 5 represents the most anomalous source while 0 is the least.
By its nature, unsupervised learning has no pre-defined metric to calculate success. However, as noted earlier, we have votes from citizen scientists for all of our light curves from BfS:MKT, which in turn have been verified by us. Throughout this work we will assume that these verified volunteer votes are an absolute ground truth against which to compare our anomaly detection methods. We can then use the positions of our citizen-science verified transients in our anomaly score-ordered list to calculate metrics of success - the more transients that are ranked as highly anomalous, the better. It is important to point out that Astronomaly never decides whether or not something is anomalous and simply orders the entire list by anomaly score. The use of information from our volunteers is what allows us to then calculate metrics of success such as recall and precision (see section 4.2). A contrasting method for cases without such citizen science data might be to define some threshold in an anomaly score-ordered list above which things are said to be anomalies, providing the binary classes on which to calculate performance.
@@ 4.1 Anomaly detection algorithms
The first of the two anomaly detection methods used herein is the Local Outlier Factor (LOF, Breuniq et al., 2000). The LOF is a density-based method that measures the deviation of the density of a given sample with respect to some user-defined number of nearest neighbours, $k$. We set $k=100$ throughout this work, approximately 1 per cent of our sample. Values of $k$ between 20 (the default) and 200 were tested, producing changes in performance at the sub per cent level. As noted earlier, without labelled data it would be difficult to optimise this (and other) hyperparameters. The LOF is local in that the resultant anomaly score depends on how isolated the object is with respect to the surrounding neighbourhood. More specifically, the LOF of datum A is a measure of the average local density of all k neighbours divided by A’s own local density. One advantage of this local approach is that it might pick up anomalies missed by global density measurements. That is, a point only a small distance from a very dense cluster might be an anomaly, whereas a point in a sparse cluster might still have the same density as its neighbours and so not be an outlier despite its low absolute density value. For more on the distinction between global and local anomalies, see the discussion in Rogers et al. (2024).
@@
4.2 Anomaly Recall and Precision
In order to assess the performance of our models we can ask out of all the anomalies that exist, what fraction did we find in the top n sources? This quantity is known as the recall (also referred to as the completeness, true positive rate or the sensitivity) and is only calculable with ground truth labels for an entire dataset, as without this it is impossible to know how many anomalies are missed (false negatives). Recall is defined as
@@ 4.2 Anomaly Recall and Precision
where for us the denominator is our total number of transients (anomalies) in the data, i.e. 168. This recall value, whilst simple, corresponds well with preferable performance consisting of many more anomalies high in the list for the user to see quickly. We can calculate the recall $R$ as a function of n until we reach the end of our dataset ($n=8874)$, by which time we will have recovered all sources in the list. We define $R_{10}$ and $R_{30}$ as the recall of transients in the top 10 and 30 per cent of our list respectively, which we will use as quantitative measures of performance throughout this work. We stress again that it is impossible to calculate this recall without labels for all transients, as without this we would not know how many transients fall below a given threshold i.e. the rate of false negatives. As such, looking at the top 10 or 30 per cent of the anomaly score-ordered lists is somewhat arbitrary and depends on, for example, the resources available for follow up or further citizen science classification.
We can also ask what fraction of the top $n$ sources are transients. This is known as the precision, also referred to in astronomical literature as the purity, defined as
@@
This can be thought of as a measure of the hit-rate of our anomalies. Namely, how many transients (anomalies) would there be in the top 10 or 30 per cent of the data? By definition, as $n$ approaches the total length of our dataset (8874 sources), the precision will tend towards the aggregate rate of transients in our data i.e. $\sim 2$ per cent. precision, unlike recall, only depends on the labels of sources above some cut-off point set by the user and so is calculable even for a largely unlabelled dataset.
@@ 4.3 Active Learning
In this work we actively label the top $\sim$ 2 per cent of each feature-model pair, which takes of order 10 minutes by an individual (AA), applying a consistent methodology throughout. That is, examples of clear variability or known transients were ranked highly ($U=4$ or 5), whilst low-significance light curves, or those showing clear false positive behaviour - e.g. where one epoch shows large changes in flux due to a different calibrator or worse image quality - were ranked at 1 or 0.
The rate of retraining in Astronomaly is entirely determined by the user(s). Any number of sources can be labelled between training iterations, with between 10 and 20 sources proving to be a qualitatively good balance between providing enough data for training whilst updating regularly enough for efficient labelling. As an example of this procedure, if the user labelled a few sources that were deemed uninteresting in the same way this would, upon retraining, down-weight further instances of similar light curves (that is, ones with similar features). However, it is worth noting that this labelling is inherently subjective and dependent on a user’s end goal. An instrument scientist interested in flux calibration would score sources differently to an astronomer interested in stellar radio flares. Active learning attempts to quantify this and balance an algorithmic definition of anomalous with a user-dependent idea about what is an interesting anomaly. Finally we note that the regressor used to model user interest is by default a RF but that using a Gaussian process provides an improvement on this for some datasets, which has been incorporated into the most recent versions of Astronomaly (Walmsley et al., 2022b; Lochner & Rudnick, 2025).
@@ 5.1 Baseline Anomaly Detection
Figure 6: Upper panel: The recall curves for both anomaly detection algorithms as applied to four groups of features derived from our radio light curves. The data are ordered according to the anomaly score from each algorithm, where for example $x=0.3$ corresponds to the most anomalous thirty per cent of the dataset. The dotted line in each corresponds to $y=x$ i.e. random chance. Lower panel: A combined plot of all eight recall lines for better comparison between feature sets. Dashed lines make comparisons (see text) for recall rates at 10 and 30 per cent of the dataset. The values in the legend correspond to the area under each recall curve (AUC).
@@
The recall curves for both anomaly detection algorithms applied to all four feature sets can be seen in Figure 6. These are normalised on both axes i.e. x=0.1 corresponds to the top 10 per cent most anomalous sources and y=0.5 corresponds to half of the bona fide 168 variables found by volunteers. Better performance corresponds to curves with higher gradient at low $n$, appearing pushed up into the upper left corner. It can be clearly seen that ($\eta,V$) show poor performance on this dataset, essentially matching random chance, in line with A23’s findings. As noted earlier, this will be due to a combination of applying a different methodology to the intended use of $\eta$ and $V$ and the inherent biases of these two statistics - e.g. $\eta$ is correlated with signal-to-noise squared (see equation 2). Furthermore, our ground truth data consist of sources chosen irrespective of their ($\eta,V$) values. Contrastingly, we see that our main feature sets are able to recover a majority of our volunteer-verified transients (anomalies) in a minority of the data. This is the first demonstration of anomaly detection as applied to radio time series.
@@
When comparing performance between anomaly detection models (i.e. for the same feature set), we see minimal variation and qualitatively similar behaviour between recall curves. For the feets and wavelet features we see ‘broken power-law’ type behaviour, with an approximately constant gradient initially, followed by a knee and then worse performance. When combining these feature sets, performance is similar, though the IF and LOF diverge more after $x=0.1$ before reconverging. Table 2 lists all $R_{10}$ and $R_{30}$ values for feature-model set pairs. The largest discrepancies between models occur at the ‘top’ of our list, with $R_{10}$ values varying by 0.1 for feets features and 0.14 for wavelets. By contrast, $R_{30}$ values differ from 0.03 to 0.07, indicating that model choice most affects the highly anomalous sources in our dataset. When combining the feets and wavelet features, this difference between $R_{10}$ and $R_{30}$ values is less pronounced. It is worth noting that several iterations of each feature-model pair were tested, with performance never varying by more than $\sim$1 per cent. As a result, all values are rounded to this precision.
@@
Differences in feature set have a much greater effect than changing the anomaly detection method - for the same model $R_{10}$ changes by factors of $\sim$3.5 between best- and worst- performing feature sets. Similarly, $R_{30}$ for the baseline parameters reaches only $\sim$0.34, whilst the wavelet features can recover up to 91 per cent of volunteer-verified transients. This underscores the importance of using features that best describe a given dataset.
@@
We can also calculate an aggregate recall performance by simply integrating the area under the curve (AUC) of each feature-model pair’s recall curve. The corresponding AUC values for each feature-model instance are listed in the legend of Figure 6. As already mentioned, differences between feature sets are greater than differences between anomaly detection models. These AUC values show that the wavelet-IF combination has the best recall performance at an AUC of 0.83. However, this is a summary over all the data and does not consider the actual finding of transients in a live setting. By this we mean that one must consider how many anomalies a human will be able to find in a reasonable amount of time. As discussed above and can be seen in the lower panel of Figure 6, if a user can only inspect the top 10 per cent of the dataset, then the feets-LOF analysis performs the best, with $R_{10}=0.52$. However, if there are resources available to inspect 30 per cent of the dataset (e.g. computation, citizen scientists, a team of experts) then the LOF performance on wavelet features recalls 91 per cent of transients, 7 per cent better than the next-best feature-model pair ($R_{30}=0.84)$.
feets Wavelets $(\eta,V)$ Combined
LOF IF LOF IF LOF IF LOF IF
$R_{10}$ 0.52, 0.52 0.42, 0.45 0.24, 0.32 0.38, 0.42 0.14, 0.16 0.12, 0.11 0.38, 0.51 0.34, 0.37
$R_{30}$ 0.74, 0.73 0.78, 0.78 0.91, 0.91 0.84, 0.84 0.34, 0.32 0.31, 0.30 0.76, 0.78 0.69, 0.74
$P_{10}$ 0.11, 0.11 0.09, 0.10 0.05, 0.06 0.07, 0.08 0.03, 0.03 0.02, 0.02 0.07, 0.10 0.07, 0.07
$P_{30}$ 0.05, 0.05 0.05, 0.05 0.06, 0.06 0.05, 0.05 0.02, 0.02 0.02, 0.02 0.05, 0.05 0.04, 0.05
@@
Table 2: Specific recall ($R$) and precision ($P$) values, rounded to two decimal places, for each feature-model pair, measured after ten and thirty per cent of the most anomalous sources have been ‘seen’. Each pair of values listed are the default, active iteration of that feature-model instance.
@@
The precision values for all feature-model pairs can be found in Table 2, again looking at 10 and 30 per cent of our anomaly score-ordered lists. For the feets and wavelet feature sets, precision is always at least a factor of 2 better than the corresponding baseline performance. In general, the same trends hold as with our recall values - differences between feature sets are greater than differences between anomaly detection models. Likewise, we can see that the feets features produce the best results when looking at the top 10 per cent of the data, with a hit-rate of over 1 in 10, compared to the baseline of 1 in 50. The results 30 per cent of the way through the anomaly score-ordered lists, as seen with the recall values, show that the wavelet-LOF instance performs the best, but only marginally. Nevertheless, we can see that for either feature set or their combination, as applied to both of our anomaly detection algorithms, the hit-rate of anomalies in the data is always a factor of 2 or greater when compared to our baseline. This directly translates into the number of sources that would need to be sifted out before finding a new transient.
@@
In summary, the choice of model used makes a small difference, though LOF is marginally preferred. When considering feature sets, the baseline features perform the worst, recovering transients almost uniformly through the anomaly score-ordered lists. The feets features perform the best at the beginning of each anomaly score-ordered list, recovering 52 per cent of transients in 10 per cent of the data and increasing the precision above random by a factor of 5. However, the wavelet features perform better when considering a larger fraction of each anomaly score-ordered list. Using the wavelet features over 90 per cent of anomalies are recovered in 30 per cent of the data, at a precision level that is three times higher than the baseline features.